# MoWE-Audio: Multitask AudioLLMs with Mixture of Weak Encoders

### v1

@@ I Introduction
In this paper, we focus on Audio Large Language Models (AudioLLMs), the class of multimodal LLMs that process and understand speech and audio inputs in conjunction with text. Existing AudioLLMs [4, 5, 6, 8, 11, 12, 13, 14, 15] are capable of executing tasks for both speech and non-speech audio inputs. Speech tasks include automatic speech recognition, speech-to-text translation and speech question answering, and non-speech audio tasks include audio captioning, sound event classification and audio question anwering. Typically, a pre-trained audio encoder is connected to a pre-trained LLM, and the model is instruction finetuned with curated audio task datasets [4, 5, 6, 7, 8]. The audio encoder is pre-trained with supervised learning [16] or self-supervised learning algorithms [17, 18, 19]. Recent AudioLLMs [4, 5, 7] have utilized the Whisper-large encoder [16], an advanced model that specializes in speech recognition and speech translation.
@@ II-A AudioLLMs
The mainstream architecture employed in multi-modal LLMs, including AudioLLMs [4, 5, 6, 7, 8] and vision-language [9, 10] models, consists (1) an encoder to embed the non-language-modality input, (2) a modality adapter and projection to process the resulting emebddings to produce tokens that align with the LLM input space, and (3) an LLM that generates free-form text based on the resulting tokens and textual instruction prompts. For instance, Qwen-Audio [5] connects a Whisper-large-v2 [16] encoder to a Qwen-7B [22] LLM. Some works use two audio encoders to capture the semantic and acoustic information in the audio inputs. SALMONN [4] connects a Whisper-large-v2 and a Beats [19] encoder to Vicuna-13B [3], and WavLLM [7] connects a Whisper-large-v2 and a WavLM-base [17] encoder to Llama-2-chat-7B [2]. We incorporate more encoders through mixture of experts routing. Furthermore, we show that even incorporating the same type of encoders can improve performance.
@@ V-A Mixture of Uniform Encoders
We perform further analysis on the design of MoWE routers in Table IV. With a single mixture, a data-dependent router achieves better performance than a data-independent router, demonstrating the effect of specialization. With two mixtures, having two data-dependent routers perform suboptimally as they may introduce excessive dynamic changes in the encoder embeddings. A combination of data-dependent router and data-indepdent router performs the best.
@@ V-C Comparison with Models with Large-Scale Training
VI Conclusion

### v4

@@ I Introduction
In this paper, we focus on Audio Large Language Models (AudioLLMs), the class of multimodal LLMs that process and understand speech and audio inputs in conjunction with text. Existing AudioLLMs [11, 12, 4, 5, 6, 8, 13, 14, 15, 16, 17] are capable of executing tasks for both speech and non-speech audio inputs. Speech tasks include automatic speech recognition, speech-to-text translation and speech question answering, and non-speech audio tasks include audio captioning, sound event classification and audio question anwering. Typically, a pre-trained audio encoder is connected to a pre-trained LLM, and the model is instruction finetuned with curated audio task datasets [11, 4, 5, 6, 7, 8]. The audio encoder is pre-trained with supervised learning [18] or self-supervised learning algorithms [19, 20, 21]. Recent AudioLLMs [11, 4, 5, 7] have utilized the Whisper-large encoder [18], an advanced model that specializes in speech recognition and speech translation.
@@ II-A AudioLLMs
The mainstream architecture employed in multi-modal LLMs, including AudioLLMs [11, 4, 5, 6, 7, 8] and vision-language [9, 10] models, consists (1) an encoder to embed the non-language-modality input, (2) a modality adapter and projection to process the resulting embeddings to produce tokens that align with the LLM input space, and (3) an LLM that generates free-form text based on the resulting tokens and textual instruction prompts. For instance, Qwen-Audio [5] connects a Whisper-large-v2 [18] encoder to a Qwen-7B [24] LLM. Some works use two audio encoders to capture the semantic and acoustic information in the audio inputs. SALMONN [4] connects a Whisper-large-v2 and a Beats [21] encoder to Vicuna-13B [3], and WavLLM [7] connects a Whisper-large-v2 and a WavLM-base [19] encoder to Llama-2-chat-7B [2]. We incorporate more encoders through mixture of experts routing. Furthermore, we show that even incorporating the same type of encoders can improve performance.
@@ V-A Mixture of Uniform Encoders
We perform further analysis on the design of MoWE routers in Table IV. With a single mixture, a data-dependent router achieves better performance than a data-independent router, demonstrating the effect of specialization. With two mixtures, having two data-dependent routers perform suboptimally as they may introduce excessive dynamic changes in the encoder embeddings. A combination of data-dependent router and data-independent router performs the best.
@@ V-C Comparison with Models with Large-Scale Training
VI Further Analysis
We evaluate MoWE on additional out-of-distribution datasets not seen at training: LibriSpeech-Other [30] for Automatic Speech Recognition, SLUE-P2-SQA [40] for Speech Question Answering, WavCaps [41] for Audio Captioning and WavCaps-AQA [41] for Audio Question Answering. The datasets are processed and the evaluation is performed with AudioBench [35]. From Table VII, we note that MoWE performs comparatively with the baseline on these out-of-distribution datasets. This is not surprising since (multimodal) LLM capabilities depend on the training data and improving generalization is still an open problem.
TABLE VII: Performance on out-of-distribution datasets not seen at training. M.J. refers to Model-as-Judge on a 0-5 grading scale.
MoWE LibriSpeech-Other SLUE-P2-SQA WavCaps WavCaps-AQA WER ()($\downarrow$)( ↓ ) M.J. ()($\uparrow$)( ↑ ) M.J. ()($\uparrow$)( ↑ ) M.J. ()($\uparrow$)( ↑ ) 5.57 3.85 1.19 1.37 5.50 3.82 1.16 1.44
We provide details on the total number of parameters, number of active parameters, and throughput measured by number of samples processed per second in Table VIII. For the count of active parameters, we assume samples are drawn from each dataset with equal probability. Throughput is computed on one Nvidia H100 GPU. Adding the mixture of weak encoders only slightly reduces throughput, as the parameter counts of the weak encoders are significantly smaller in comparison with the LLM (9-95M vs. 8B).
TABLE VIII: Computational efficiency at inference time.
MoWE # total parameters # active parameters # samples per second 8.73B 8.73B 1.92 8.96B 8.91B 1.83
VII Conclusion
@@
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.